### v2

# SAGE: Strategy-Adaptive Generation Engine for Query Rewriting

Abstract
Query rewriting is pivotal for enhancing dense retrieval, yet current methods demand large-scale supervised data or suffer from inefficient reinforcement learning (RL) exploration. In this work, we first establish that guiding Large Language Models (LLMs) with a concise set of expert-crafted strategies substantially improves retrieval effectiveness on challenging benchmarks, including HotpotQA, FEVER, NFCorpus, and SciFact. Building on this insight, we introduce the Strategy-Adaptive Generation Engine (SAGE), which operationalizes these strategies in an RL framework. SAGE introduces two novel reward shaping mechanisms- Strategic Credit Shaping (SCS) and Contrastive Reward Shaping (CRS)-to deliver more informative learning signals. This strategy-guided approach not only achieves new state-of-the-art NDCG@10 results, but also uncovers a compelling emergent behavior: the agent learns to select optimal strategies, reduces unnecessary exploration, and generates concise rewrites, lowering inference cost without sacrificing performance. Our findings demonstrate that strategy-guided RL, enhanced with nuanced reward shaping, offers a scalable, efficient, and more interpretable paradigm for developing the next generation of robust information retrieval systems.
SAGE: Strategy-Adaptive Generation Engine for Query Rewriting
Teng Wang OPPO Research Institution wt0318@connect.hku.hk          Hailei Gong Tsinghua University gonghl18@mails.tsinghua.edu.cn
Changwang Zhang OPPO Research Institution zhangchangwang@oppo.com          Jun Wang OPPO Research Institution wangjun7@oppo.com
1 Introduction
Effective information retrieval (IR) is increasingly reliant on dense retrieval systems, which map queries and documents into a shared semantic space. The performance of these systems, however, is fundamentally bound by the quality of the input query. To bridge the significant gap between a user’s initial intent and a query optimized for machine comprehension, query rewriting has emerged as a critical component. While LLMs have shown significant promise for data generation, comprehension, and reasoning Achiam et al. (2023); Wang et al. (2025a); Yang et al. (2025b, a); Wang et al. (2025b), current methodologies face two primary obstacles: traditional supervised fine-tuning demands large-scale, costly manual annotations, whereas modern RL approaches, such as PPO Schulman et al. (2017) and GRPO Shao et al. (2024), often struggle with inefficient exploration. This inefficiency not only hampers the discovery of optimal rewriting strategies but can also result in unstable training dynamics and even catastrophic failures, where models produce incoherent or irrelevant outputs.
While prior work by Li et al. (2024) has explored strategy-based prompting, we find their strategies, specifically designed for sparse retrieval in general web search, struggle to generalize to the nuanced demands of dense retrieval and exhibit limited effectiveness on specialized benchmarks. To address these challenges, we propose five novel query-rewriting strategies specifically designed for dense retrieval scenarios. These strategies significantly enhance the ability of LLMs to effectively reformulate queries, consistently outperforming previous method Li et al. (2024) on diverse and challenging benchmarks, including HotpotQA Yang et al. (2018), FEVER Thorne et al. (2018), NFCorpus Boteva et al. (2016), and SciFact Wadden et al. (2020).
Figure 1: An overview of the SAGE framework. SAGE operationalizes expert-crafted strategies within a reinforcement learning loop. The policy model generates a strategy-guided rewrite, which is evaluated against the environment to produce an initial reward (NDCG@10). This reward is then transformed by our novel shaping modules-Strategic Credit Shaping or Contrastive Reward Shaping-to create a potent learning signal for the agent.
The successful application of on-policy RL to fine-tune LLMs is pioneered by algorithms such as PPO Schulman et al. (2017), which in turn inspired a family of variants including GRPO Shao et al. (2024), VAPO Yue et al. (2025), and DAPO Yu et al. (2025). While these powerful algorithms have enabled new capabilities in query rewriting for LLMs, many prior approaches Jiang et al. (2025) simply adopt them as generic optimization tools, without incorporating domain-specific guidance tailored to the query rewriting task.
To bridge this gap, we introduce the Strategy-Adaptive Generation Engine (SAGE), as demonstrated in Fig. 1, a framework that directly integrates our human-designed strategies into the GRPO algorithm, steering LLMs toward more effective query-rewriting policies. To further refine the learning signal, we propose two novel reward shaping schemes. In addition to using the direct NDCG@10 score, we introduce Strategic Credit Shaping (SCS), which assigns credit based on the average performance of each strategy, and Contrastive Reward Shaping (CRS), which transforms absolute scores into relative performance measures.
By reproducing prior approaches Li et al. (2024); Jiang et al. (2025), we identify a significant form of reward hacking Weng (2024); Wang et al. (2024a) inherent in the query rewriting task. When fine-tuned on datasets such as HotpotQA Yang et al. (2018), models tend to converge to the trivial policy of directly copying the input query provided in the prompt, as modern retrievers like BGE-en-v1.5 Xiao et al. (2023) already achieve strong baseline performance with the original queries. Notably, even semantically appropriate rewrites frequently result in lower retrieval scores, discouraging exploration and trapping the agent in a local optimum that precludes the discovery of potentially superior strategies.
To address this issue, we promote exploration by modifying the prompt and penalizing outputs identical to the original query. This approach encourages the agent to venture beyond the safe default. The effectiveness of this strategy is further validated by detailed ablation studies in Section 4.3.
We evaluate SAGE on two challenging datasets: HotpotQA Yang et al. (2018) and NFCorpus Boteva et al. (2016). Our results demonstrate that SAGE achieves state-of-the-art retrieval effectiveness as measured by NDCG@10. Notably, we observe an emergent behavior where SAGE learns a more efficient reasoning process, substantially reducing token usage. This improved efficiency directly translates to lower inference latency and reduced computational costs.
Our main contributions are summarized as follows:
  1. 1.
    We show that prompting LLMs with a small set of interpretable strategies substantially improves query rewriting quality, establishing the performance upper bound attainable through prompting alone.
  2. 2.
    We introduce SAGE, a novel RL framework that systematically integrates explicit strategies into the learning process. SAGE autonomously adapts strategy selection via our proposed reward-shaping mechanisms, enabling more effective policy optimization.
  3. 3.
    We establish a new state-of-the-art in dense retrieval effectiveness (NDCG@10) using SAGE, while identifying a notable emergent behavior: SAGE learns a more efficient reasoning process, substantially reducing inference latency and computational costs.
  4. 4.
    We provide comprehensive analyses and ablation studies highlighting the critical importance of forced exploration. Our results underscore the necessity of an explicit penalty mechanism to avoid reward hacking, offering valuable insights for effectively training RL-based rewriting models.
2 Methodology
The core idea of our methodology is to replace black-box RL optimization with explicit, human-interpretable decision-making. To this end, we propose the Strategy-Adaptive Generation Engine (SAGE), which reformulates query rewriting as a structured, strategy-driven decision process. Our framework directly incorporates expert-crafted strategies into RL training to systematically guide both exploration and policy learning, as illustrated in Figure 1.
2.1 Problem Formulation
We frame query rewriting as an RL task. Given an initial user query $q_{orig}$ and a document collection $\mathcal{D}$, the objective is to learn a policy $\pi$ that generates an improved query $q$. The environment evaluates the rewritten query using NDCG@10, yielding an initial reward $r_{\text{orig}}=\text{NDCG@10}(q,\mathcal{D})$, which is further refined by our reward shaping methods (see Section 2.4) to produce the final reward $r_{\text{final}}$. The agent thus aims to maximize the expected final reward: $\mathbb{E}_{q\sim\pi(\cdot|q_{orig})}[r_{\text{final}}]$. We use GRPO Shao et al. (2024) as our optimization algorithm and incorporate a set of expert-crafted rewriting strategies $\mathcal{S}=\{s_{1},s_{2},\dots,s_{5}\}$, providing explicit guidance for query reformulation.
2.2 Explicit Strategic Primitives
To move beyond generic prompting, we introduce five explicit, expert-crafted query rewriting strategies, each tailored to address specific challenges in dense retrieval. These strategies systematically alter the semantic structure of queries to mitigate common retrieval failures. The primitives, summarized in Table 4, are not mutually exclusive and often involve inherent trade-offs (e.g., semantic expansion may improve recall at the expense of precision). Collectively, they constitute the discrete action space from which our SAGE framework dynamically selects and applies. Detailed prompt templates for each strategy are provided in Appendix A.1.
2.3 The SAGE Framework
SAGE operationalizes our expert-crafted strategies by embedding them within an on-policy RL algorithm. In contrast to black-box methods, SAGE transforms the task from unconstrained text generation into a structured, two-part action selection process, making the agent’s decision-making more explicit and interpretable. The overall workflow is illustrated in Figure 1.
For each input query, the policy model generates a batch of $N$ rollouts, where each action is a structured pair, $\{q_{i},s_{i}\}$, comprising the rewritten query and the integer ID of the selected strategy. This formulation requires the LLM to produce not only an effective rewrite but also an explicit strategy choice, rendering its intent interpretable.
Each rewritten query $q_{i}$ is evaluated by the retrieval environment to obtain an initial reward $r_{\text{orig},i}$, measured by NDCG@10. In conventional RL, this raw score serves as the final reward; however, we argue that it provides insufficient learning signal, as it neither credits effective strategy selection nor clearly distinguishes superior rewrites within a batch. To address this, we process the initial rewards with a dedicated reward shaping module (see Section 2.4), which transforms the raw scores into a more informative signal $r_{\text{final},i}$ for policy updates.
Figure 2: An illustration of our two proposed reward shaping mechanisms. (a) SCS solves the credit assignment problem by grouping rollouts based on their chosen strategy, ranking these strategies by their average performance, and then scaling the individual rewards by their strategy’s rank. (b) CRS sharpens the learning signal by normalizing each reward against a baseline (e.g., the batch median), reframing the objective as outperforming the typical performance.
2.4 Reward Shaping Mechanisms
To provide the agent with a more informative learning signal, we introduce two novel reward shaping mechanisms that go beyond simply using the raw NDCG@10 score. While the direct score is a standard choice, it provides only limited feedback: it fails to explicitly reward the selection of high-performing strategies and does not encourage intra-batch competition, as it treats all positive outcomes similarly-making little distinction between adequate and exceptional rewrites. To address these shortcomings, we propose two new mechanisms that transform the raw score into a more effective learning objective, as illustrated in Figure 2.
Strategic Credit Shaping (SCS):
This method aims to explicitly solve the credit assignment problem by rewarding the agent for selecting high-performing strategies. Within a given batch of rollouts, we first group them by their chosen strategy $s_{i}$. We then compute the average initial reward, $\bar{r}_{\text{orig}}(s_{i})$, for each strategy group and rank them based on this score. The final reward for an individual rollout is its original score scaled by the inverse of its strategy’s rank:
$r_{\text{SCS},i}=\frac{r_{\text{orig},i}}{\text{rank}(s_{i})}$ (1)
This mechanism directly encourages the agent to converge on strategies that are collectively more effective, providing a clearer signal for strategic decision-making.
Contrastive Reward Shaping (CRS):
This method introduces intra-batch competition by normalizing rewards against a dynamic baseline, effectively penalizing underperforming rewrites while rewarding those that surpass the batch’s typical performance. The final reward is the advantage over this baseline:
$r_{\text{CRS},i}=r_{\text{orig},i}-\text{baseline}$ (2)
This forces the agent to learn policies that outperform the batch’s typical performance, rather than simply achieving any positive score. By creating a clearer distinction between superior and mediocre rewrites, CRS sharpens the reward landscape and accelerates learning.
2.5 Countering Reward Hacking with Forced Exploration
Reproducing previous work Jiang et al. (2025); Li et al. (2024), we observe pervasive reward hacking in query rewriting: agents exploit reward function loopholes to achieve high scores without fulfilling the intended objective. In our setting, the strong baseline performance of modern retrievers allows agents to obtain high rewards simply by copying the original query, a strategy the policy model quickly adopts since it has direct access to the input. As a result, the agent becomes trapped in a deceptive local optimum, stifling exploration of potentially superior rewriting strategies. To directly counteract this, we introduce two mechanisms designed to force exploration.
Exploration Penalty.
We apply a simple penalty term to disincentivize the agent from reverting to the trivial policy of copying the input. The final reward, $r_{\text{final}}$, is calculated by applying this penalty to the base reward signal, $r_{\text{base}}$, which represents the output from our main reward calculation (either direct NDCG@10, SCS, or CRS). The relationship is formally defined as:
$r_{\text{final}}=\begin{cases}r_{\text{base}}-p&\text{if }q_{orig}=q\\ r_{\text{base}}&\text{otherwise}\end{cases}$ (3)
where $q_{orig}$ is the original query, $q$ is the rewritten query, and $p$ is a fixed penalty hyperparameter. This formulation ensures that the agent is explicitly penalized only when it outputs a query identical to the input, thereby directly encouraging the exploration of novel rewrites.
Figure 3: Performance scaling laws for query rewriting using the BGE-base-en-v1.5 retriever across four benchmark datasets: HotpotQA, FEVER, NFCorpus, and SciFact. Each subplot compares the upper-bound performance (NDCG@10) of our Strategy-Guided Prompting (SGP) against the strategies from DMQR-RAGLi et al. (2024), evaluated via a Best-of-N methodology across the Qwen3 model series Yang et al. (2025a).
Proactive Exploration Prompting.
In contrast to prior work Li et al. (2024), which instructs the model to keep the query unchanged if a better one is not found, our prompting philosophy is designed to actively encourage the agent to explore alternative rewrites.
We validate the effectiveness of these forced exploration mechanisms in our ablation studies in Section 4.3.
Method HotpotQA NFCorpus
NDCG@10 $\uparrow$ Avg. Tokens $\downarrow$ NDCG@10 $\uparrow$ Avg. Tokens $\downarrow$
Baselines
   Original Query 0.6633 0 0.3677 0
   Qwen3-4B 0.6366 1598 0.3527 911
   Qwen3-8B 0.4295 1966 0.3298 885
   Qwen3-14B 0.5251 1682 0.3473 586
   GPT-4.1 0.7118 297 0.3711 205
   GPT-o4-mini 0.6915 776 0.3809 440
   Claude-Sonnet-4-thinking 0.6515 1451 0.3701 1152
   Gemini-2.5-Flash 0.6425 1614 0.3689 957
   Gemini-2.5-Pro 0.6671 1986 0.3663 1096
   Deepseek-R1 0.6262 2182 0.3270 1251
   DeepRetrieval + Qwen3-4B 0.6681 232 0.3676 343
   DMQR-RAG + Qwen3-4B 0.5812 798 0.1808 1032
Our Method (SAGE)
   SAGE (Direct) 0.6894 92 0.3776 229
   SAGE-SCS 0.6955 66 0.3967 139
   SAGE-CRS 0.6918 69 0.4035 154
Table 1: SAGE achieves state-of-the-art performance with remarkable efficiency. This table compares our SAGE framework, fine-tuned on Qwen3-4B, against strong baselines including the re-evaluated DeepRetrieval method and significantly larger proprietary models like Gemini-2.5-Pro and GPT-o4-mini. All models are evaluated with a maximum response length of 4096 tokens. The results demonstrate that SAGE not only outperforms other RL approaches but also achieves retrieval effectiveness (NDCG@10) competitive with these massive models, while requiring substantially fewer tokens.
3 Experiment
3.1 Effectiveness of Rewriting Strategies
To empirically validate our five expert-crafted strategies (detailed in Section 2.2 and Appendix A.1), we conduct a comprehensive scaling law analysis using BGE-en-base-v1.5 Xiao et al. (2023) as the retriever. We compare our Strategy-Guided Prompting (SGP) to the strategies in DMQR-RAG Li et al. (2024) under a Best-of-N methodology, with results shown in Figure 3. Across all four challenging datasets, including HotpotQA Yang et al. (2018), FEVER Thorne et al. (2018), NFCorpus Boteva et al. (2016), and SciFact Wadden et al. (2020), our SGP consistently and substantially outperforms the baseline. This demonstrates that our strategies provide a more effective upper-bound on performance by successfully guiding the LLM towards higher-quality query rewrites. Similar trends of SGP outperforming the baseline are observed when using the Contriever Izacard et al. (2021) retriever, with detailed results provided in Appendix A.2.
3.2 Comparison with State-of-the-Art Baselines
Having established the upper-bound potential of our expert-crafted strategies (Section 3.1), we now evaluate the full SAGE framework. Our preliminary analysis indicates that for datasets like FEVER and SciFact, the potential for improvement via RL is constrained, due to either a high performance baseline from the original query or a high exploration cost required for marginal gains. Consequently, to provide a more meaningful evaluation of SAGE’s optimization capabilities, we focus our main RL experiments on HotpotQA and NFCorpus, which present a larger performance gap and thus a more dynamic learning environment.
Unless otherwise specified, all experiments are conducted using BGE-en-base-v1.5. Detailed training settings are provided in Appendix A.3.
We benchmark SAGE against a diverse suite of strong baselines to demonstrate its effectiveness. This includes state-of-the-art open-source models Qwen3-4B, Qwen3-8B, Qwen3-14B, GPT-4.1, GPT-o4-mini, Claude-Sonnect-4, Gemini-2.5-Flash, Gemini-2.5-Pro, Deepseek-R1 Yang et al. (2025a); Team et al. (2023); Achiam et al. (2023); Anthropic (2025); Guo et al. (2025). For a fair comparison under our controlled experimental setup, we also re-evaluate prior RL-based methods, including DeepRetrieval Jiang et al. (2025) and the strategies from DMQR-RAG Li et al. (2024). Against these baselines, we report the performance of SAGE enhanced with our two novel reward shaping mechanisms, SAGE-SCS and SAGE-CRS.
The results presented in Table 1 clearly demonstrate the dual advantage of our SAGE framework. In terms of retrieval effectiveness, our SAGE variants establish a new state-of-the-art. Notably, SAGE, fine-tuned on Qwen3-4B, consistently outperforms specialized RL-based approaches like DeepRetrieval also fine-tuned from Qwen3-4B in our experiment setting. More strikingly, our method achieves performance that is competitive with, and in some cases superior to, significantly larger and more powerful proprietary models such as Gemini-2.5-Pro, GPT-o4-mini.
More striking than the performance gains is the emergent efficiency of our SAGE framework. We observe that SAGE consistently generates rewrites using substantially fewer tokens than competing methods, a phenomenon that is even more pronounced when using our SCS and CRS reward shaping schemes. We attribute this to the agent learning to prioritize more direct reasoning paths, effectively discovering a shortcut to high-quality rewrites without being explicitly optimized for brevity.
This finding has profound practical implications. In a production environment, the long, static portion of our strategic prompt can be pre-processed and its KV cache stored. At inference time, the system only needs to generate a short reasoning, significantly reducing latency and computational cost. This architectural advantage, which decouples the pre-computable prompt context from the short dynamic generation, makes SAGE a highly attractive solution for real-world, performance-critical retrieval systems.
Figure 4: Qualitative comparison illustrating the mechanism behind SAGE’s dual advantage. The baseline model (left), lacking strategic guidance, engages in a verbose and convoluted reasoning process, resulting in a suboptimal query. In contrast, SAGE (right) leverages an explicit strategy to find a direct and efficient path to a more semantically accurate rewrite, achieving superior retrieval effectiveness with a fraction of the token generation cost.
Method HotpotQA NFCorpus
NDCG@10 $\uparrow$ Avg. Tokens $\downarrow$ NDCG@10 $\uparrow$ Avg. Tokens $\downarrow$
Baseline (No Strategic Guidance)
   DeepRetrieval 0.6681 232 0.3676 343
Our Method (with Strategic Guidance)
   SAGE (Direct) 0.6894 92 0.3775 229
   SAGE-SCS 0.6955 66 0.3966 139
   SAGE-CRS 0.6918 69 0.4035 154
Table 2: Ablation study on the impact of strategic guidance. We compare the performance of a standard RL baseline against our SAGE framework, which incorporates expert-crafted strategies. The results highlight the significant gains in both effectiveness (NDCG@10) and efficiency (Avg. Tokens) brought by our strategic guidance.
4 Ablation Study
To isolate and quantify the impact of our core contributions, we conduct a series of ablation studies. Our experiments are designed to answer several key questions: (1) What is the performance gain from our expert-crafted strategies compared to an unguided RL baseline? And what is the contribution of our novel reward shaping mechanisms (SCS and CRS) compared to a direct reward? (2) How critical is the exploration penalty for overcoming reward hacking?
4.1 The Impact of Strategic Guidance
The most significant finding from our ablation study is the dramatic impact of strategic guidance. As shown in Table 2, removing our expert-crafted strategies and reverting to a "black-box" RL approach results in a significant degradation in performance. While this change leads to a notable drop in NDCG@10 across both datasets, the more striking effect is on efficiency, where the average response length increases dramatically. This demonstrates that models trained without strategic guidance are not only less effective but also substantially less efficient.
To provide a more intuitive understanding, Fig.  4 presents a direct case comparison. The output from the model trained without strategies is generic and fails to capture the query’s nuance, whereas the output from SAGE is scientifically precise and far more effective for retrieval.
Furthermore, we see the additional benefits of our reward shaping mechanisms. On both datasets, SAGE-SCS and SAGE-CRS further improve upon the SAGE (Direct) baseline, pushing the performance ceiling even higher while maintaining superior efficiency. This confirms that while strategic guidance is the foundational improvement, our novel reward shaping schemes provide a crucial secondary optimization.
Experimental Condition NDCG@10 $\uparrow$ Modification Rate $\uparrow$
Conservative Prompt ("keep unchanged") 0.671 0.509
Proactive Exploration Prompt 0.694 0.951
Proactive Prompt + Penalty ($p=0.02$) 0.692 0.998
Table 3: Ablation study on different exploration mechanisms for SAGE-SCS on the HotpotQA dataset. We compare three settings: (1) using a conservative prompt, (2) using a proactive exploration prompt, and (3) combining the proactive prompt with an exploration penalty.
4.2 Analysis of Training Dynamics
Beyond final performance, the training dynamics also reveal the superiority of the SAGE framework. As illustrated in Appendix A.4, the baseline model trained without strategies quickly gets trapped in a local optimum. Its performance stagnates after only a few training steps, and its response length remains high and volatile throughout the process.
In stark contrast, SAGE demonstrates continuous performance improvement during training. More importantly, we observe a compelling emergent behavior: as SAGE learns to better utilize its strategies, it also learns to achieve the optimal rewrite with a more concise reasoning process. This results in a steady and significant decrease in the average response length over time, proving that our framework not only learns more effectively but also more efficiently.
4.3 The Critical Role of the Exploration Penalty
In reproducing prior work Jiang et al. (2025); Li et al. (2024), we identify a significant challenge in this task: the agent’s strong tendency to revert to a trivial policy of outputting the original query without modification. We attribute this reward-hacking behavior to two primary causes. First, the high baseline performance of modern retrievers makes the "do-nothing" action a safe, high-reward option. This issue is exacerbated by instructions in prior work like Li et al. (2024), which explicitly instructs the agent to preserve the original query if a rewrite is deemed unnecessary, reinforcing this conservative policy from the start of training. Second, many powerful retrievers have been trained on these benchmarks, making them brittle and highly sensitive to phrasing, where even minor, semantically-sound modifications can lead to a sharp drop in performance. This combination traps the agent in a deceptive local optimum, making any exploration a high-risk endeavor.
To break this cycle and force meaningful exploration, we implement a two-pronged approach. First, in contrast to prior work Li et al. (2024) which explicitly instructs the agent to keep the original query if no better one is found, our prompt philosophy actively encourages exploration, detailed is demonstrated in Appendix A.1. Second, and more critically, we introduce a simple penalty term, subtracting a fixed value p=0.05 from the reward whenever the generated query is identical to the original, directly disincentivizing this reward-hacking behavior.
To validate the effectiveness of these mechanisms, we conducted an ablation study using the SAGE-SCS model on HotpotQA. The results, shown in Table 3, clearly demonstrate the value of our approach: simply replacing the conservative prompt with our proactive exploration prompt substantially increases both NDCG@10 and the modification rate, highlighting the critical role of encouraging exploration. However, the use of an exploration penalty reveals a critical trade-off. While it drives the modification rate to 0.998, we observe a slight decrease in NDCG@10. This suggests that in cases where the original query is already optimal, enforcing modifications via a penalty may slightly impair peak performance. Thus, although the penalty is highly effective for maximizing exploration and mitigating reward hacking, its application should be carefully balanced against the specific task’s need for exhaustive exploration versus optimal end performance.
5 Conclusion
In this work, we introduced SAGE, a novel framework designed to address the black-box nature and inefficient exploration of conventional RL-based query rewriting. By integrating expert-crafted strategies and novel reward shaping mechanisms (SCS and CRS) into the RL loop, SAGE achieves state-of-the-art retrieval effectiveness and emergent efficiency, substantially reducing inference costs. Importantly, our approach effectively mitigates reward hacking by explicitly promoting exploration and penalizing trivial, unmodified rewrites. These findings demonstrate that strategy-guided RL not only enhances effectiveness and efficiency but also leads to more transparent and controllable retrieval systems.
6 Limitation
6.1 Evaluation Datasets
One of the key innovations introduced in this paper is the query rewrite strategy or policy. This approach relies on manual analysis to craft rewriting strategies tailored to specific datasets. However, this methodology poses certain limitations. In particular, it is less applicable to domains such as sparse retrieval and SQL-based retrieval, where the query structure is already well-defined and leaves little room for rewriting strategies. Consequently, the benefits of our approach are constrained in these scenarios. Our work is therefore primarily focused on the dense retrieval setting, where more flexibility in query reformulation exists.
Appendix A Appendix
A.1 Strategy-Guided Prompting
As referenced in Section 2.2, we present detailed strategies and the complete prompt used to guide our policy model in Table 4 and Figure 5.
Figure 5: Full Prompt Structure for Strategy-Guided Rewriting. The figure presents the complete instructional prompt provided to the LLM agent. It specifies the overall task, formally defines each of our five expert-crafted strategies, and provides illustrative examples to guide their application.
A.2 SGP Performance with Contriever
To confirm that our findings are robust and not specific to a single retriever, we replicated the experiment using Contriever Izacard et al. (2021). As detailed in Fig 6, our strategy again demonstrates a significant performance margin over the methods proposed by DMQR-RAGLi et al. (2024).
Figure 6: Performance scaling laws for query rewriting using the Contriever Izacard et al. (2021) retriever across four benchmark datasets: HotpotQA Yang et al. (2018), FEVER Thorne et al. (2018), NFCorpus Boteva et al. (2016), and SciFact Wadden et al. (2020). Each subplot compares the upper-bound performance (NDCG@10) of our SGP against the strategies from Li et al. (2024), evaluated via a Best-of-N methodology across the Qwen3 model series Yang et al. (2025a). The results demonstrate the superior scaling potential and higher performance ceiling of our approach.
A.3 Training Hyperparameters
For all experiments, we fine-tune the Qwen3-4B Yang et al. (2025a) model using Verl Sheng et al. (2024) and GROP Shao et al. (2024), a reinforcement learning framework. We experiment with rollout numbers of 16 and 32. All training is conducted on 8 NVIDIA A100 GPUs. Unless otherwise specified, we use the default hyperparameters provided by Verl.
A.4 Detailed Training and Response Length Curves
This section provides the detailed learning curves that support our analysis of training dynamics in the main text. Fig. 7 and Fig. 8 illustrate the validation NDCG@10 and the average training rollout response length as a function of training steps. These plots compare the learning trajectories of our SAGE variants against a baseline model trained without strategies. The results demonstrate that while the baseline model’s performance quickly stagnates, our SAGE framework exhibits a superior learning trajectory, achieving continuous improvement in effectiveness while simultaneously learning to generate more concise responses.
Figure 7: Training dynamics on the HotpotQA dataset. Each row corresponds to a different method: (from top to bottom) baseline without strategies, SAGE (Direct), SAGE-CRS, and SAGE-SCS. The left column plots the validation NDCG@10 (evaluated every 30 steps) over training steps, while the right column plots the average response length. The baseline stagnates early, whereas all SAGE variants show continuous improvement in NDCG@10 and a steady decrease in response length.
Figure 8: Training dynamics on the NFCorpus dataset. Each row corresponds to a different method. The left column plots validation NDCG@10(evaluated every 30 steps), and the right column shows average response length. As with HotpotQA, the baseline model’s performance plateaus while exhibiting high response length. In contrast, our SAGE variants, particularly SAGE-SCS and SAGE-CRS, demonstrate both stable performance gains and a significant reduction in response length over the course of training.
A.5 Related Work
The Evolution of Query Rewriting
Query rewriting has long been a cornerstone of IR, aimed at bridging the semantic gap between user intent and document representations. Early approaches typically relies on rule-based methods, thesaurus expansion, or statistical machine translation techniques. Although effective in specific contexts, these approaches often lack robustness and required extensive domain-specific feature engineering or sizable parallel corpora Rocchio Jr (1971); Zhai and Lafferty (2001); Abdul-Jaleel et al. (2004); Dalton et al. (2014); Xu et al. (2009); Xiong and Callan (2015). The advent of pre-trained sequence-to-sequence models Liu et al. (2019); Devlin et al. (2018), such as BART Lewis et al. (2020) and T5 Raffel et al. (2020), significantly reshape query rewriting, framing it explicitly as a supervised fine-tuning task. However, this new paradigm introduces a critical bottleneck: heavy reliance on large-scale, high-quality annotated query pairs, which are costly and labor-intensive to construct.
Reinforcement Learning for Query Rewriting
To alleviate reliance on explicit supervision, recent approaches leverage RL to guide the generative capabilities of LLMs using weaker signals Schulman et al. (2017); Shao et al. (2024); Jiang et al. (2025); Li et al. (2024); Wang et al. (2025b, d, c). However, the application of these powerful algorithms often reveals significant limitations. A common approach is to treat the LLM as a black box, applying on-policy algorithms like PPO without modification Jiang et al. (2025), which overlooks the need for specialized guidance. Other lines of work focus on prompt engineering, but strategies tailored for sparse retrieval in general web search Li et al. (2024) exhibit limited effectiveness when transferred to dense retrieval tasks. Furthermore, even approaches that integrate reward models, such as Wang et al. (2025d), can be flawed; they often rely on an arbitrary, static fusion of scores from multiple fine-tuned reward models, lacking dynamic feedback from the actual retrieval environment. This highlights a fundamental gap: existing RL-based methods frequently neglect the nuanced, domain-specific guidance required to fully exploit the potential of dense retrieval systems. In contrast, our SAGE framework addresses this gap by introducing a novel layer of explicit, human-designed strategic guidance. It achieves this through two core innovations: first, by equipping the agent with a set of fine-grained strategies for query augmentation, and second, by proposing novel reward calculation schemes like SCS and CRS to provide a more nuanced and effective learning signal.
Challenge in RL: Reward Hacking
Beyond the lack of strategic guidance, applying RL to generative tasks introduces a further challenge. A pervasive issue in reinforcement learning is reward hacking, where an agent learns to maximize a misspecified or ambiguous reward function in unintended ways. This often results in the agent securing high scores by exploiting loopholes, rather than mastering the desired behavior or accomplishing the true underlying objective Weng (2024); Wang et al. (2024a); Wen et al. (2025); Wang et al. (2024b).
In the context of query rewriting, this problem manifests in a particularly deceptive form. Since the policy model has access to the original query, it can discover a trivial, low-effort strategy: leave the query unchanged. Given that modern retrievers can achieve high baseline performance on the original queries for many benchmarks, this "do nothing" policy is often reinforced with a strong, positive reward. This effectively traps the agent in a local optimum, allowing it to "hack" the reward by avoiding the risk of exploration altogether. Our work directly confronts this challenge by introducing mechanisms to force meaningful exploration.